# Sigma: Semantically Informative Pre-training for Skeleton-based Sign Language Understanding (Under review)

### v1

@@ H.1 Optimising local cluster-wise contrastive learning strategies
Figure 9: Illustration of the local cluster-wise contrastive learning process inspired by Chen et al. (2020); Radford et al. (2021); Li et al. (2022a); Hou et al. (2024). The similarity matrix is computed between visual tokens and textual clusters. For each row, the maximum similarity value is obtained using a row max operation. The resulting scores are passed through a softmax function to normalise the scores. Finally, in-batch local cluster contrastive learning is applied to pull semantically aligned visual-text pairs (highlighted in yellow) closer, while pushing apart unaligned pairs. This process enables localised semantic grounding by focusing on the most relevant visual-text associations within each cluster.
H.1.1 Row-wise operations
@@
H.1.2 Local-level scoring
@@
H.2 Should the text encoder be trainable in Sigma’s pre-training?
@@
GloFE-VN Lin et al. (2023) 21.06 12.34 8.68 6.68 21.37 21.56 12.74 9.05 7.06 21.75
Conv-GRU Camgoz et al. (2018) 16.72 8.95 6.31 4.82 16.25 16.11 8.85 6.18 4.58 16.10
I3D-transformer Shi et al. (2022) 18.26 10.26 7.17 5.60 18.88 18.31 10.15 7.19 5.56 18.64
OpenASL Shi et al. (2022) 20.10 11.81 8.43 6.57 20.43 20.92 12.08 8.59 6.72 21.02
Uni-Sign Li et al. (2025) 50.84 37.82 29.83 24.16 44.58 49.35 36.32 28.55 23.14 43.22
$C^{2}RL$ Chen et al. (2025) - - - - - 31.46 21.85 16.58 13.21 31.36
@@
GloFE-VN Lin et al. (2023) 14.90 7.30 3.90 2.20 12.60
YouTube-ASL Uthus et al. (2023) 37.80 24.10 16.90 12.40 -
MSLU Zhou et al. (2024) 20.10 7.70 - 2.40 17.20
SLT-IV Tarrés et al. (2023) 34.00 19.30 12.20 8.00 -
$C^{2}RL$ Chen et al. (2025) 29.10 18.60 12.90 9.40 27.00
FLa-LLM Chen et al. (2024) 29.80 19.00 13.30 9.70 27.80
@@
SLRT Camgoz et al. (2020) 37.47 24.67 16.86 11.88 37.96 37.38 24.36 16.55 11.79 36.74
ConSLT Fu et al. (2023) - - - 14.80 41.46 - - - 14.53 40.98
SignBT Zhou et al. (2021a) 51.46 37.23 27.51 20.80 49.49 51.42 37.26 27.76 21.34 49.31
MMTLB Chen et al. (2022a) 53.81 40.84 31.29 24.42 53.38 53.31 40.41 30.87 23.92 53.25
SLTUNET Zhang et al. (2023) - - - 23.99 53.58 54.98 41.44 31.84 25.01 54.08
TS-SLT Chen et al. (2022b) 55.21 42.31 32.71 25.76 55.10 55.44 42.59 32.87 25.79 55.72
CV-SLT Zhao et al. (2024a) 58.05 44.73 35.14 28.24 56.36 58.29 45.15 35.77 28.94 57.06
@@
SLRT Camgoz et al. (2020) 21.03 9.97 5.96 4.04 20.51 20.00 9.11 4.93 3.03 19.67
GASLT Yin et al. (2023) - - - - - 19.90 9.94 5.98 4.07 20.35
MSLU Zhou et al. (2024) 33.28 21.31 - 10.27 33.13 33.97 22.20 - 11.42 33.8
NSLT Camgoz et al. (2018) 34.22 19.72 12.24 7.96 34.28 34.16 19.57 11.84 7.56 34.54
GFSLT-VLP Zhou et al. (2023) 39.20 25.02 16.35 11.07 36.70 39.37 24.93 16.26 11.00 36.44
FLa-LLM Chen et al. (2024) - - - - - 37.13 25.12 18.38 14.20 37.25
$C^{2}RL$ Chen et al. (2025) - - - - - 49.32 36.28 27.54 21.61 48.21
Uni-Sign Li et al. (2025) 55.30 42.21 32.94 26.25 56.03 55.08 42.14 32.98 26.36 56.51
SignLLM Gong et al. (2024) 42.45 26.88 17.90 12.23 39.18 39.55 28.13 20.07 15.75 39.91
Sign2GPT Wong et al. (2024) - - - - - 41.75 28.73 20.60 15.40 42.36
@@
H.3 Cooperate different modalities during pre-training
@@
Appendix I Complete results of experiments
@@
Appendix K Visualising cross-modal alignment with t-SNE
To visually evaluate the quality of the learned representations, we randomly sample eighty paired sign sequences and their corresponding translated text(s) to visualise the 2D distribution of sign and text features using t-distributed stochastic neighbour embedding (t-SNE). This dimensionality reduction technique preserves local neighbourhood structure in high-dimensional data, making it especially effective for examining the alignment between modalities in the latent space. The features used in this analysis are derived from the downstream model of Sigma, ensuring that they reflect task-specific representations. Specifically, sign features are extracted from the output of the sign encoder, while text features are obtained immediately after the text embedding layer of the SGT decoder. This design choice reflects the decoding process during inference, where the sign encoder outputs are fed into the SGT decoder to generate textual representations. For each task, we compare the features learned by models trained from scratch to those initialised with pre-trained parameters from the pre-training stage. This comparison provides insight into how pre-training shapes cross-modal alignment and reveals the effectiveness of Sigma in learning semantically grounded representations.
@@
Appendix L Ethics Statement
@@
Appendix M Reproducibility Statement
@@
Appendix N Use of Large Language Models

### v2

@@ I.1 Optimising local cluster-wise contrastive learning strategies
Figure 9: Illustration of the computation of our local sign-to-text cluster-wise similarity inspired by (chen2020simple; radford2021learning; li2022blip; hou2024bagformer). The similarity matrix $M$ is computed between the sign tokens of each sample and all textual clusters. For each row, the maximum similarity value is obtained using a row max operation. The resulting values are passed through a softmax-weighted sum function to obtain the local similarity scores. Finally, in-batch local cluster contrastive learning is applied to pull semantically aligned visual-text pairs (highlighted by light yellow borders) closer, while pushing apart unaligned pairs. This process enables localised semantic grounding by focusing on the most relevant visual-text associations within each cluster.
I.1.1 Row-wise operations
@@
I.1.2 Local-level scoring
@@
I.2 Should the text encoder be trainable in Sigma’s pre-training?
@@
GloFE-VN (lin2023gloss) 21.06 12.34 8.68 6.68 21.37 21.56 12.74 9.05 7.06 21.75
Conv-GRU (camgoz2018neural) 16.72 8.95 6.31 4.82 16.25 16.11 8.85 6.18 4.58 16.10
I3D-transformer (shi2022open) 18.26 10.26 7.17 5.60 18.88 18.31 10.15 7.19 5.56 18.64
OpenASL (shi2022open) 20.10 11.81 8.43 6.57 20.43 20.92 12.08 8.59 6.72 21.02
Uni-Sign (Li2025sign) 50.84 37.82 29.83 24.16 44.58 49.35 36.32 28.55 23.14 43.22
$C^{2}RL$ (chen2025c) - - - - - 31.46 21.85 16.58 13.21 31.36
@@
GloFE-VN (lin2023gloss) 14.90 7.30 3.90 2.20 12.60
YouTube-ASL (uthus2023youtube) 37.80 24.10 16.90 12.40 -
MSLU (zhou2024scaling) 20.10 7.70 - 2.40 17.20
SLT-IV (tarres2023sign) 34.00 19.30 12.20 8.00 -
$C^{2}RL$ (chen2025c) 29.10 18.60 12.90 9.40 27.00
FLa-LLM (chen2024factorized) 29.80 19.00 13.30 9.70 27.80
@@
SLRT (camgoz2020sign) 37.47 24.67 16.86 11.88 37.96 37.38 24.36 16.55 11.79 36.74
ConSLT (fu2023token) - - - 14.80 41.46 - - - 14.53 40.98
SignBT (zhou2021improving) 51.46 37.23 27.51 20.80 49.49 51.42 37.26 27.76 21.34 49.31
MMTLB (chen2022simple) 53.81 40.84 31.29 24.42 53.38 53.31 40.41 30.87 23.92 53.25
SLTUNET (zhang2023sltunet) - - - 23.99 53.58 54.98 41.44 31.84 25.01 54.08
TS-SLT (chen2022two) 55.21 42.31 32.71 25.76 55.10 55.44 42.59 32.87 25.79 55.72
CV-SLT (zhao2024conditional) 58.05 44.73 35.14 28.24 56.36 58.29 45.15 35.77 28.94 57.06
@@
SLRT (camgoz2020sign) 21.03 9.97 5.96 4.04 20.51 20.00 9.11 4.93 3.03 19.67
GASLT (yin2023gloss) - - - - - 19.90 9.94 5.98 4.07 20.35
MSLU (zhou2024scaling) 33.28 21.31 - 10.27 33.13 33.97 22.20 - 11.42 33.8
NSLT (camgoz2018neural) 34.22 19.72 12.24 7.96 34.28 34.16 19.57 11.84 7.56 34.54
GFSLT-VLP (zhou2023gloss) 39.20 25.02 16.35 11.07 36.70 39.37 24.93 16.26 11.00 36.44
FLa-LLM (chen2024factorized) - - - - - 37.13 25.12 18.38 14.20 37.25
$C^{2}RL$ (chen2025c) - - - - - 49.32 36.28 27.54 21.61 48.21
Uni-Sign (Li2025sign) 55.30 42.21 32.94 26.25 56.03 55.08 42.14 32.98 26.36 56.51
SignLLM (gong2024llms) 42.45 26.88 17.90 12.23 39.18 39.55 28.13 20.07 15.75 39.91
Sign2GPT (wong2024sign2gpt) - - - - - 41.75 28.73 20.60 15.40 42.36
@@
I.3 Cooperate different modalities during pre-training
@@
Appendix J Complete results of experiments
@@
Appendix L Visualising cross-modal alignment with t-SNE
To visually evaluate the quality of the learned representations, we randomly sample eighty paired sign sequences and their corresponding translated text(s) to visualise the 2D distribution of sign and text features using t-distributed stochastic neighbour embedding (t-SNE). This dimensionality reduction technique preserves local neighbourhood structure in high-dimensional data, making it especially effective for examining the alignment between modalities in the latent space. The features used in this analysis are derived from the downstream model of Sigma, ensuring that they reflect task-specific representations. Specifically, sign features are extracted from the output of the sign encoder, while text features are obtained immediately after the text embedding layer of the SGT decoder. This design choice reflects the decoding process during inference, where the sign encoder outputs are fed into the SGT decoder to generate textual representations. For each task, We compare the features learned by models initialised with only the mT5 weight to those initialised with pre-trained parameters from the pre-training stage. This comparison provides insight into how pre-training shapes cross-modal alignment and reveals the effectiveness of Sigma in learning semantically grounded representations.
@@
Appendix M Additional ablation study
I In the main paper, we assess the impact of each core component by varying the number of fusion layers, balancing local-global feature modelling ($\alpha$), and adjusting the weights of text matching and language modelling ($\beta$). In this section, we extend those analyses with additional experiments targeting extreme configurations. Together with the t-SNE visualisations, these results provide further evidence that each component plays a valuable role in bridging the cross-modality gap. They also highlight that the effectiveness of pre-training depends on balanced feature modelling and structured modality interaction, which collectively ensure semantically aligned SLU representations across ISLR, CSLR, and SLT tasks.
Table 24: Impact of SignEF.
Layers WLASL2000 CSL-Daily
ISLR CSLR SLT
P-I↑ P-C↑ WER↓ B@1↑ B@4↑ R@L↑
0 62.82 60.38 26.86 56.43 26.80 56.01
1 63.17 60.77 26.58 56.79 27.50 56.64
2 63.79 61.40 26.12 56.83 28.24 58.04
3 64.22 62.09 26.53 56.72 27.93 57.36
4 63.97 61.70 26.40 56.80 27.98 57.86
5 64.40 62.32 26.69 56.79 28.09 57.56
When the number of fusion layers is set to zero (see Table 24), the model disables the SignEF mechanism, leading to a performance drop across all SLU tasks. For instance, ISLR on the WLASL2000 dataset records its lowest performance in this setting, with per-instance and per-class accuracies of 62.82% and 60.38%, respectively. Similarly, CSLR on CSL-Daily yields a higher WER of 26.86%. SLT performance also declines, with BLEU-1 at 56.43, BLEU-4 at 26.80, and ROUGE-L at 56.01, indicating a broader degradation across tasks. These results indicate the importance of SignEF in differentiating visually similar gestures, especially in tasks like ISLR and SLT where semantic precision and expressive generation are crucial, which beyond what WER alone can capture. The absence of early visual-linguistic interaction hampers the ability of Sigma to establish strong alignment between sign and text representations. In contrast, even minimal integration of SignEF yields consistent gains, emphasising the importance of early-stage modality fusion for deeper semantic grounding and improved sequence modelling.
Table 25: Local-global feature balancing.
Alpha WLASL2000 CSL-Daily
ISLR CSLR SLT
P-I↑ P-C↑ WER↓ B@1↑ B@4↑ R@L↑
0.0 63.85 61.72 27.12 56.00 27.16 57.32
0.2 64.14 62.14 26.52 56.11 27.99 57.89
0.4 64.30 62.24 26.55 56.65 28.12 57.82
0.5 64.40 62.32 26.12 56.83 28.24 58.04
0.6 64.14 62.09 27.05 56.82 27.99 58.10
0.8 64.14 62.08 26.65 56.21 27.66 57.65
1.0 63.67 61.55 27.27 55.99 27.00 57.24
Beta WLASL2000 CSL-Daily
ISLR CSLR SLT
P-I↑ P-C↑ WER↓ B@1↑ B@4↑ R@L↑
0.0 64.07 61.82 26.77 56.01 27.63 57.36
0.2 64.30 62.21 26.62 56.31 27.82 57.88
0.4 64.40 62.15 26.43 56.38 27.88 57.76
0.5 64.27 62.03 26.12 56.83 28.24 58.04
0.6 64.40 62.32 26.33 56.00 27.79 57.43
0.8 64.12 62.22 26.27 55.58 27.72 57.36
1.0 63.92 61.76 26.55 55.26 26.98 56.88
The performance at the two extremes of the $\alpha$ parameter (0.0 and 1.0) highlights the importance of balancing local and global feature alignment (see Table 25). When $\alpha$ is set to 0.0, the model relies solely on global alignment, ignoring fine-grained local interactions. This leads to degraded performance across all SLU tasks. On the other hand, when $\alpha$ is set to 1.0, the model depends entirely on local contrastive learning, without leveraging global class-level alignment. Comparing these extremes reveals that ISLR and SLT performance suffers more when relying exclusively on local features, while CSLR shows slightly better results under local-only alignment. These findings suggest that neither local nor global alignment alone is sufficient to capture the full spectrum of semantic relationships required for SLU. The strongest performance is achieved when both are combined, reinforcing the necessity of joint local-global modelling for learning semantically meaningful and generalizable representations across diverse SLU tasks.
The results for $\beta=0.0$ and $\beta=1.0$ in Table 26 illustrate the importance of balancing text matching and language modelling during training. At $\beta=0.0$, when the objective is driven entirely by text matching, performance drops across all SLU tasks. At the other end, with $\beta=1.0$, the model relies only on language modelling, yielding lower SLT results than at $\beta=0.0$, but more pronounced degradation in ISLR and similar degree of degradation in CSLR. These observations show that placing too much emphasis on either modality-specific alignment or generative fluency fails to deliver consistent performance across tasks. The performance drop at both ends highlights the need for a balanced approach, where integrating text matching and language modelling allows the model to align cross-modal semantics while producing coherent textual outputs.
The extent of performance degradation under extreme parameter settings in our proposed method varies across tasks and evaluation metrics. Overall, the most substantial decline occurs when the SignEF module is removed, underscoring its essential role in enabling effective cross-modal interaction. One exception is the BLEU-1 score, where the decrease is relatively modest, suggesting that surface-level lexical matching may be less dependent on early fusion compared to deeper semantic metrics like BLEU-4 or ROUGE-L. When comparing the impact of imbalanced feature modelling versus unbalanced text matching and language modelling, the former tends to result in more pronounced performance drops. An exception occurs when $\beta=1.0$, where BLEU-4 reaches its second lowest value, indicating that overreliance on language modelling can impair the ability of the model to preserve fine-grained sign-text correspondence. These findings highlight the need for balanced integration across both feature learning and text objectives to achieve robust and semantically aligned SLU performance.
Appendix N Ethics Statement
@@
Appendix O Reproducibility Statement
@@
Appendix P Use of Large Language Models